Solve the IVP:

y" + 3y' + 2y = 6((delta)base 6)(t)

with conditions: y(0) = 0, y'(0) =-1.
1
Expert's answer
2020-09-21T12:41:23-0400

Given differential equation is

"y''+3y'+2y = \delta_6(t)" with conditions is "y(0)=0, y'(0)=-1" .

Taking Laplace transform on both sides, we get

"L(y'')+ 3L(y')+2L(y )= L(\delta_6(t)) \\
\implies s^2 Y + 1 + 3 sY + 2 Y = e^{-6s}" where "Y = L(y)" .

"\implies (s^2+3s+2) Y = e^{-6s} -1 \\
\implies Y =\frac{ e^{-6s} -1}{s^2+3s+2} \\
\implies y =L^{-1}(\frac{ e^{-6s} -1}{s^2+3s+2})"

So, "y =\left(\left(\left(- e^{12 - t} + e^{6}\right) \delta_6(t )- 1\right) e^{t} + 1\right) e^{- 2 t}"